Question is that I have x+y=5
x+z=6
y+z=7
Please solve this question.
I can't do this.
Regards
Faraz Saleem

Answer by rothauserc(4718)

we are given
1) x + y = 5
2) x + z = 6
3) y + z = 7
subtract equation 1 from equation 2, we get
z - y = 1
then add equation 3 to the above result
2z = 8
z = 4
from equation 2 we have
x + 4 = 6
x = 2
from equation 3 we have
y = 3
solution set is x = 2, y = 3, z = 4